OpenStudy (anonymous):

help1 question this is your medal

OpenStudy (anonymous):

How can polynomial functions be written when given the zeros?

OpenStudy (anonymous):

If the zeroes of the polynomial are \[ r_1, r_2, \cdots, r_n, \](with repeats allowed), then the polynomials possible are \[ a \cdot \left( x - r_1 \right) \cdot \left( x - r_2 \right) \cdots \left( x - r_n \right), \]for any real constant a.

OpenStudy (anonymous):

thank you kind sir

OpenStudy (anonymous):

You're welcome!